The kinetic energy of a particle executing shm is 16 j when it is in its mean position If the amplitude of oscillation is 25 and the mass of particle is 5.12kg the time period of its oscillation is

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Pallavi

x=Xsinwt

v=dx/dt =Xwcoswt

at mean position particle has only kinetic energy

at mean position velocity is V=Xw

K.E=1/2 *m*V2

  16= 1/2 m X2w2

 32=5.12 (25)2 (2∏/T)2

now you can calculate T



Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed

solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

shanky shanky
14 years ago

i am agree with previous  sol.

so ans.=52.34sec

ts tabse n
5 years ago
Am kam kam kam kam ka. Ka. Ka. Kam kam kam kam kam. Kam. Oma mao mo amo maom kam a ka Oma moam kam palm kam kam pa kam Kamchatka Oma Oma Oma ago. As. Ago. Ago mao mao mao moam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam kam oam aim oam oam oam oam ka. O mao mA. O a oao a of oaxoa ka ka ka ka ka ka ka oao ao a ka ka ka oao ao ao Alcoa ao ao ao AOC a ka oa